• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 18:15:30 on June17,2010

    Entry number 327101

    keyword=cool-down progress and cryo ioc problem

    Target cool-down progressing well: after cooled down the line, I started opening J-T to cool the target. The supply/return from ESR
    are all at 20s (22/26)K while our target supply and return are
    still high (134/146)K. Scattering chamber vacuum went up when J-T first opened (warm gas coming first due to line warmed before cold
    gas arrived). Now vacuum is pretty good at 2-3*10-6.
    We lost EPIC readback values from ESR. Talked to MCC and I was told dthat probably it's due to the cryo ioc problem and the Guides should already called the cryo on-call to look at the problem.